OpenStudy (anonymous):

Please help 1. 10.6 < b (1 point) * –18 * –9 * 7 * 14 2. m greater then 7/12 (1 point) * 1 * –1 * –9 * –5 3. 8 < x(7 – x) (1 point) * 2 * 8 * –1 * 0 12. The function b(n) = 6n represents the number of light bulbs b(n) that are needed for n chandeliers. How many lightbulbs are needed for 15 chandeliers? (1 point) * 90 lightbulbs * 2 lightbulbs * 96 lightbulbs * 80 lightbulbs For questions 13–14, tell whether the sequence is arithmetic. If it is, what is the common differenc

OpenStudy (anonymous):

Bamboo plants grow rapidly. A bamboo plant is 130 inches tall. Tomorrow it will be 143 inches tall, the next day it will be 156 inches tall, and on the next day it will be 169 inches tall. Write a rule to represent the height of the bamboo plant as an arithmetic sequence. How tall will the plant be in 13 days? (1 point) * A(n) = 130 + (n – 1)13; 286 inches * A(n) = 130 + (n – 1)13; 299 inches * A(n) = 130 + 13n; 286 inches * A(n) = 130 + 13n; 299 inches